# Scaf-GRPO: Scaffolded Group Relative Policy Optimization for Enhancing LLM Reasoning

### v1

@@ Abstract
Figure 1: Scaf-GRPO overcomes the learning cliff with minimal guidance, outperforming vanilla GRPO grpo and the prefix-based LUFFY yan2025learning across challenging math benchmarks on Qwen2.5-Math-7B. By injecting strategic, hierarchical hints, our method unlocks the model’s potential on difficult problems, achieving superior overall performance.
@@ 1  Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in complex reasoning tasks across diverse domains such as mathematics, programming, and logic DeepSeekR1 ; openaio1 ; muennighoff2025s1 ; min2024imitate . A key driver of these advancements is Reinforcement Learning from Verifier Rewards (RLVR) DeepSeekR1 ; simplerlzoo ; liu2025understandingr1zeroliketrainingcritical , a paradigm where models learn to generate sophisticated reasoning paths by exploring diverse strategies and receiving feedback on their final outcomes. This approach eliminates the need for expensive, step-by-step human annotations by rewarding only the final correct answer, enabling models to autonomously discover effective problem-solving procedures DeepSeekR1 .
However, the efficacy of RLVR is severely constrained by a fundamental challenge we formalize as the “learning cliff.” This phenomenon occurs when a model confronts a subset of problems that lie significantly beyond its current capabilities. For these problems, all exploratory attempts consistently fail, leading to two critical and cascading consequences: (1) Reward Signal Loss: The model receives a persistent zero-reward signal for this entire subset of problems. (2) Vanishing Gradients: In algorithms like GRPO grpo , the advantage signal provides the learning gradient. When all rewards are zero, the advantage collapses to zero, providing no gradient for the policy to learn from DAPO .
@@
To address the learning cliff, a prevailing strategy has emerged: incorporating off-policy guidance from a more capable “teacher” policy yan2025learning ; Prefix-RFT ; StepHint ; BREAD . These methods typically work by providing the student model with a prefix of a correct “golden” solution and tasking it with generating the remainder of the reasoning path. While this ensures a positive reward signal, this prefix-continuation paradigm introduces significant issues. It creates a distributional mismatch between the teacher-generated prefix and the student-generated suffix, necessitating complex algorithmic corrections like policy shaping yan2025learning or hybrid SFT-RL objectives Prefix-RFT that can introduce bias and training instability. More critically, this “on-rails” guidance forces the model down a predetermined path, stifling its ability to explore alternative, potentially more novel or efficient, reasoning strategies.
To address this challenge, we propose Scaf-GRPO (Scaffolded Group Relative Policy Optimization). Our framework is inspired by the pedagogical theory of Scaffolding Scaffolding , a teaching method of providing temporary support that fades as learners improve. We apply this principle by providing hierarchical, minimal, progressive assistance to help the model bridge its capability gaps, rather than enforcing a rigid solution prefix. This in-prompt scaffolding approach is guided by two primary objectives: first, to maintain policy consistency by having the model process both the problem and the hint under a single, unified policy, thereby avoiding the distributional mismatches of prefix-based methods. Second, to preserve exploration flexibility, as our hints act as “signposts” rather than “railroads,” guiding the model without fixing its path and allowing it to discover its own unique solution strategies.
@@ 2  Related Work
The success of DeepSeek-R1 DeepSeekR1 establishes Reinforcement Learning from Verifier Reward (RLVR) as a paradigm for enhancing the reasoning capabilities of Large Language Models (LLMs). In RLVR, models are trained using feedback from an external verifier that provides an outcome-based reward (e.g., correct/incorrect) for a generated solution. The success of DeepSeek-R1 DeepSeekR1 demonstrates that even with sparse, binary rewards, models can learn reasoning strategies. Subsequent research has built upon this foundation, focusing on enhancing algorithmic stability through debiasing techniques liu2025understandingr1zeroliketrainingcritical ; DAPO , or designing more informative rewards to improve sample efficiency, such as using length penalties to mitigate overthinking L1 ; jin2025recut ; chen2024not or token-level signals to provide denser feedback High-Entropy-token ; Dual-token .
@@
To overcome the learning cliff, a phenomenon where a persistent lack of positive rewards renders difficult problems invisible to the learning gradient DAPO , researchers incorporate guidance from a “teacher” policy. The prevailing strategy is to provide the student model with a prefix of a “golden” trajectory and task it with generating the continuation yan2025learning ; Prefix-RFT ; StepHint ; BREAD ; ma2025learning . Different methods introduce variations on this theme. For instance, Luffy yan2025learning mixes a complete expert trajectory with multiple model-generated rollouts in one batch. Prefix-RFT Prefix-RFT employs a cosine decay schedule to adjust the length of the guiding prefix. More recently, StepHint StepHint provides multi-level hints of varying lengths, allowing the model to explore from multiple starting points. However, this prefix-continuation paradigm introduces challenges. It breaks policy consistency by mixing trajectories from two different distributions, which necessitates complex algorithmic patches yan2025learning ; Prefix-RFT ; StepHint . Furthermore, forcing the model down a predetermined path stifles exploration, limiting its ability to discover novel reasoning strategies. Our work provides effective guidance while circumventing these issues.
@@ 3.1  Preliminaries: Group Relative Policy Optimization (GRPO)
Group Relative Policy Optimization (GRPO) grpo is an on-policy RL algorithm for training LLMs that eliminates the need for a trainable value function. For a given prompt $q$, the policy $\pi_{\theta}$ generates a group of $N$ trajectories, $\mathcal{G}=\{o_{1},\dots,o_{N}\}$. After obtaining a terminal reward $R(o_{i})$ for each trajectory from a verifier, GRPO computes a normalized advantage $\hat{A}_{i}$ as: $\hat{A}_{i}=\frac{R(o_{i})-\mu_{\mathcal{G}}}{\sigma_{\mathcal{G}}+\epsilon_{\text{std}}}$, where $\mu_{\mathcal{G}}$ and $\sigma_{\mathcal{G}}$ are the mean and standard deviation of rewards in the group $\mathcal{G}$, and $\epsilon_{\text{std}}$ is a small constant for numerical stability. The policy is then updated by maximizing a clipped surrogate objective:
@@ 3.2  The Scaf-GRPO Framework
Figure 4: Prompt for hint injection.
To provide the minimal necessary guidance, the framework executes a deterministic search through this hierarchy, proceeding from the most abstract to the most concrete hint ($H_{\text{knowledge}}\rightarrow H_{\text{planning}}\rightarrow H_{\text{solution}}$). Within each tier, guidance is offered incrementally. The search terminates as soon as the model generates a correct solution, thereby identifying the minimal effective guidance required. A detailed description of this progressive exploration algorithm is provided in Appendix C.1.
@@
In the second case, the learning cliff scenario ($\forall o_{i}\in\mathcal{G},R(o_{i})=0$), standard GRPO fails. The uniform zero rewards cause the advantage calculation to collapse ($\mu_{\mathcal{G}}=0,\sigma_{\mathcal{G}}=0$), leading to a null advantage $\hat{A}_{i}=0$ and a vanishing policy gradient. Here, Scaf-GRPO intervenes by constructing the augmented batch $\mathcal{G}_{\text{final}}$. This restores the gradient by ensuring $\mu_{\mathcal{G}_{\text{final}}}>0$, which in turn guarantees a non-zero advantage signal $\hat{A}_{i}^{\prime}$. Critically, this intervention preserves the on-policy principle. Unlike off-policy methods that import trajectories from a different policy $\pi_{\phi}$ and require high-variance importance sampling corrections (e.g., using a ratio $\frac{\pi_{\theta}}{\pi_{\phi}}$), the guided trajectory $o_{h}^{*}$ is sampled directly from the current policy $\pi_{\theta}$. The probability ratio is therefore a standard on-policy ratio computed on a modified input, which is inherently more stable. This targeted, on-policy intervention transforms an unproductive, zero-gradient sample into a valuable learning opportunity without compromising the integrity of the optimization process.
@@
   SimpleRL-Zero simplerlzoo 23.3 13.3 55.0 31.6 76.8 37.2 60.8 42.6
   Oat-Zero liu2025there 30.0 16.7 62.5 34.6 78.0 41.0 62.9 46.5
   LUFFY yan2025learning 33.3 16.7 62.5 33.8 75.2 41.7 62.7 46.6
@@ 4.1  Experimental Setups
Our training data is derived from the DeepScaleR-Preview-Dataset deepscaler2025 . We employ a dynamic filtering strategy that aligns the dataset with each model’s initial capabilities. Based on preliminary evaluation, we classify problems as “Too Easy” (discarded), “Too Hard” (retained), or “Potentially Solvable” (50% subsampled). This curates a challenging yet tractable training set focused on the frontier of the model’s abilities. For this dataset, we generate our three-tiered hints by prompting the DeepSeek-R1 model DeepSeekR1 with ground-truth solution steps. Further details on our data filtering strategy and the hint generation process are provided in Appendix A.1 and Appendix A.2, respectively.
@@
We benchmark Scaf-GRPO against three distinct classes of baselines: (1) Vanilla GRPO grpo , the standard on-policy algorithm without guidance. This serves as our baseline to quantify the gains from our scaffolding mechanism. (2) Leading GRPO implementations, including Simple-RL simplerlzoo and Oat-Zero liu2025there , to contextualize our performance against highly-optimized public benchmarks. (3) LUFFY yan2025learning , a representative of RLVR with off-policy guidance. This provides a direct comparison between the dominant prefix-continuation strategy and our in-prompt scaffolding approach.
@@
We evaluate on diverse mathematics benchmarks, including GaoKao2023en chinesegaokao2024gaokao2023en , AIME24 aime24 , AIME25 aime25 , AMC amc23 , MATH-500 hendrycks2021math , and OlympiadBench he2024olympiadbench . To assess out-of-distribution (OOD) generalization, we evaluate on the scientific reasoning benchmark, GPQA-Diamond rein2024gpqa . For all benchmarks, we report pass@1 accuracy via greedy decoding. Vanilla GRPO is trained with our data and hyperparameters, and LUFFY on our data with its original parameters. For Simple-RL and Oat-Zero, we evaluate their publicly available weights.
@@
We train all models for 10 epochs using the verl framework sheng2025hybridflow , reporting results from the best-performing checkpoint. The maximum response length is 2048 tokens (8192 for the LongCoT model). Consistent with recent studies liu2025understandingr1zeroliketrainingcritical ; DAPO ; yan2025learning , we set the KL divergence penalty to zero to maximize policy exploration. A comprehensive list of hyperparameters is detailed in Appendix B.
@@ 4.2  Main Results
To verify that the benefits of Scaf-GRPO are not confined to a single model family, we extend our evaluation to the Llama-3.2-3B-Instruct model dubey2024llama . As detailed in Table 1, our framework demonstrates strong generalization. While vanilla GRPO provides a significant uplift over the base model, Scaf-GRPO achieves a further relative improvement of 10.3% in average performance. This confirms Scaf-GRPO is a model-agnostic method, capable of enhancing reasoning abilities beyond the Qwen series.
@@ 4.3  Ablation Study
We conduct a series of ablation studies on the Qwen2.5-Math-7B model (see Table 2).
Necessity of the guidance exemption period.
To validate our guidance exemption period (Phase 1), we tested a variant applying scaffolding from the start of training. As detailed in Appendix F, this resulted in a 9.2% relative performance drop for the Qwen2.5-Math-7B model compared to our full framework. This confirms that an initial autonomous learning period is crucial to prevent over-reliance on hints and foster more robust, independent reasoning.
@@
Our methodology is founded on the hypothesis that progressive guidance, from abstract concepts to concrete steps, is superior to simply providing a direct solution. To test this, we evaluate a “Solution-Only” variant that bypasses the hierarchy and immediately provides the most concrete hint. This results in a significant performance degradation of 4.9% compared to the full model. This confirms our hypothesis: compelling the model to first engage with higher-level reasoning fosters more generalizable skills.

### v2

@@ Abstract
Figure 1: Scaf-GRPO overcomes the learning cliff with minimal guidance, outperforming vanilla GRPO (Shao et al., 2024) and the prefix-based LUFFY (Yan et al., 2025) across challenging math benchmarks on Qwen2.5-Math-7B. By injecting strategic, hierarchical hints, our method unlocks the model’s potential on difficult problems, achieving superior overall performance.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in complex reasoning tasks across diverse domains such as mathematics, programming, and logic (Guo et al., 2025; Jaech et al., 2024; Muennighoff et al., 2025; Min et al., 2024). A key driver of these advancements is Reinforcement Learning from Verifier Rewards (RLVR) (Guo et al., 2025; Zeng et al., 2025; Liu et al., 2025b), a paradigm where models learn to generate sophisticated reasoning paths by exploring diverse strategies and receiving feedback on their final outcomes. This approach eliminates the need for expensive, step-by-step human annotations by rewarding only the final correct answer, enabling models to autonomously discover effective problem-solving procedures (Guo et al., 2025).
However, the efficacy of RLVR is severely constrained by a fundamental challenge we formalize as the “learning cliff.” This phenomenon occurs when a model confronts a subset of problems that lie significantly beyond its current capabilities. For these problems, all exploratory attempts consistently fail, leading to two critical and cascading consequences: (1) Reward Signal Loss: The model receives a persistent zero-reward signal for this entire subset of problems. (2) Vanishing Gradients: In algorithms like GRPO (Shao et al., 2024), the advantage signal provides the learning gradient. When all rewards are zero, the advantage collapses to zero, providing no gradient for the policy to learn from (Yu et al., 2025).
@@
To address the learning cliff, a prevailing strategy has emerged: incorporating off-policy guidance from a more capable “teacher” policy (Yan et al., 2025; Huang et al., 2025; Zhang et al., 2025a; b). These methods typically work by providing the student model with a prefix of a correct “golden” solution and tasking it with generating the remainder of the reasoning path. While this ensures a positive reward signal, this prefix-continuation paradigm introduces significant issues. It creates a distributional mismatch between the teacher-generated prefix and the student-generated suffix, necessitating complex algorithmic corrections like policy shaping (Yan et al., 2025) or hybrid SFT-RL objectives (Huang et al., 2025) that can introduce bias and training instability. More critically, this “on-rails” guidance forces the model down a predetermined path, stifling its ability to explore alternative, potentially more novel or efficient, reasoning strategies.
To address this challenge, we propose Scaf-GRPO (Scaffolded Group Relative Policy Optimization). Our framework is inspired by the pedagogical theory of Scaffolding (Berk and Winsler, 1995), a teaching method of providing temporary support that fades as learners improve. We apply this principle by providing hierarchical, minimal, progressive assistance to help the model bridge its capability gaps, rather than enforcing a rigid solution prefix. This in-prompt scaffolding approach is guided by two primary objectives: first, to maintain policy consistency by having the model process both the problem and the hint under a single, unified policy, thereby avoiding the distributional mismatches of prefix-based methods. Second, to preserve exploration flexibility, as our hints act as “signposts” rather than “railroads,” guiding the model without fixing its path and allowing it to discover its own unique solution strategies.
@@ 2 Related Work
The success of DeepSeek-R1 (Guo et al., 2025) establishes Reinforcement Learning from Verifier Reward (RLVR) as a paradigm for enhancing the reasoning capabilities of Large Language Models (LLMs). In RLVR, models are trained using feedback from an external verifier that provides an outcome-based reward (e.g., correct/incorrect) for a generated solution. The success of DeepSeek-R1 (Guo et al., 2025) demonstrates that even with sparse, binary rewards, models can learn reasoning strategies. Subsequent research has built upon this foundation, focusing on enhancing algorithmic stability through debiasing techniques (Liu et al., 2025b; Yu et al., 2025), or designing more informative rewards to improve sample efficiency, such as using length penalties to mitigate overthinking (Aggarwal and Welleck, 2025; Jin et al., 2025; Chen et al., 2024) or token-level signals to provide denser feedback (Wang et al., 2025b; a).
@@
To overcome the learning cliff, a phenomenon where a persistent lack of positive rewards renders difficult problems invisible to the learning gradient (Yu et al., 2025), researchers incorporate guidance from a “teacher” policy. The prevailing strategy is to provide the student model with a prefix of a “golden” trajectory and task it with generating the continuation (Yan et al., 2025; Huang et al., 2025; Zhang et al., 2025a; b; Ma et al., 2025). Different methods introduce variations on this theme. For instance, Yan et al. (2025) mix a complete expert trajectory with multiple model-generated rollouts in one batch. Huang et al. (2025) employ a cosine decay schedule to adjust the length of the guiding prefix. More recently, Zhang et al. (2025a) provide multi-level hints of varying lengths, allowing the model to explore from multiple starting points. However, this prefix-continuation paradigm introduces challenges. It breaks policy consistency by mixing trajectories from two different distributions, which necessitates complex algorithmic patches (Yan et al., 2025; Huang et al., 2025; Zhang et al., 2025a). Furthermore, forcing the model down a predetermined path stifles exploration, limiting its ability to discover novel reasoning strategies. Our work provides effective guidance while circumventing these issues.
@@ 3.1 Preliminaries: Group Relative Policy Optimization (GRPO)
Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is an on-policy RL algorithm for training LLMs that eliminates the need for a trainable value function. For a given prompt $q$, the policy $\pi_{\theta}$ generates a group of $N$ trajectories, $\mathcal{G}=\{o_{1},\dots,o_{N}\}$. After obtaining a terminal reward $R(o_{i})$ for each trajectory from a verifier, GRPO computes a normalized advantage $\hat{A}_{i}$ as: $\hat{A}_{i}=\frac{R(o_{i})-\mu_{\mathcal{G}}}{\sigma_{\mathcal{G}}+\epsilon_{\text{std}}}$, where $\mu_{\mathcal{G}}$ and $\sigma_{\mathcal{G}}$ are the mean and standard deviation of rewards in the group $\mathcal{G}$, and $\epsilon_{\text{std}}$ is a small constant for numerical stability. The policy is then updated by maximizing a clipped surrogate objective:
@@ 3.2 The Scaf-GRPO Framework
To provide the minimal necessary guidance, the framework executes a deterministic search through this hierarchy, proceeding from the most abstract to the most concrete hint ($H_{\text{knowledge}}\rightarrow H_{\text{planning}}\rightarrow H_{\text{solution}}$). Within each tier, guidance is offered incrementally. The search terminates as soon as the model generates a correct solution, thereby identifying the minimal effective guidance required. A detailed description of this progressive exploration algorithm is provided in Appendix D.1.
@@
In the second case, the learning cliff scenario ($\forall o_{i}\in\mathcal{G},R(o_{i})=0$), standard GRPO fails. The uniform zero rewards cause the advantage calculation to collapse ($\mu_{\mathcal{G}}=0,\sigma_{\mathcal{G}}=0$), leading to a null advantage $\hat{A}_{i}=0$ and a vanishing policy gradient. Here, Scaf-GRPO intervenes by constructing the augmented batch $\mathcal{G}_{\text{final}}$. This restores the gradient by ensuring $\mu_{\mathcal{G}_{\text{final}}}>0$, which in turn guarantees a non-zero advantage signal $\hat{A}_{i}^{\prime}$. Critically, this intervention preserves the on-policy principle. Unlike off-policy methods that import trajectories from a different policy $\pi_{\phi}$ and require high-variance importance sampling corrections (e.g., using a ratio $\frac{\pi_{\theta}}{\pi_{\phi}}$), the guided trajectory $o_{h}^{*}$ is sampled directly from the current policy $\pi_{\theta}$. The probability ratio is therefore a standard on-policy ratio computed on a modified input, which is inherently more stable. We explicitly avoid the distributional mismatch of off-policy alternatives defined by ratios like $\frac{\pi_{\theta}(\cdot|q)}{\pi_{\theta_{\text{old}}}(\cdot|q\oplus h^{*})}$, which are shown to destabilize training in Appendix F.5. Instead, by conditioning both the current and old policies on the identical hint-augmented prompt, Scaf-GRPO ensures a stable learning signal. This targeted, on-policy intervention transforms an unproductive, zero-gradient sample into a valuable learning opportunity without compromising the integrity of the optimization process.
@@
   SimpleRL-Zero Zeng et al. (2025) 23.3 13.3 55.0 31.6 76.8 37.2 60.8 42.6
   Oat-Zero Liu et al. (2025a) 30.0 16.7 62.5 34.6 78.0 41.0 62.9 46.5
   LUFFY Yan et al. (2025) 33.3 16.7 62.5 33.8 75.2 41.7 62.7 46.6
@@ 4.1 Experimental Setups
Our training data is derived from the DeepScaleR-Preview-Dataset (Luo et al., 2025). We employ a dynamic filtering strategy that aligns the dataset with each model’s initial capabilities. Based on preliminary evaluation, we classify problems as “Too Easy” (discarded), “Too Hard” (retained), or “Potentially Solvable” (50% subsampled). This curates a challenging yet tractable training set focused on the frontier of the model’s abilities. For this dataset, we generate our three-tiered hints by prompting the DeepSeek-R1 model (Guo et al., 2025) with ground-truth solution steps. Further details on our data filtering strategy and the hint generation process are provided in Appendix B.1 and Appendix B.2, respectively.
@@
We benchmark Scaf-GRPO against three distinct classes of baselines: (1) Vanilla GRPO (Shao et al., 2024), the standard on-policy algorithm without guidance. This serves as our baseline to quantify the gains from our scaffolding mechanism. (2) Leading GRPO implementations, including Simple-RL (Zeng et al., 2025) and Oat-Zero (Liu et al., 2025a), to contextualize our performance against highly-optimized public benchmarks. (3) LUFFY (Yan et al., 2025), a representative of RLVR with off-policy guidance. This provides a direct comparison between the dominant prefix-continuation strategy and our in-prompt scaffolding approach.
@@
We evaluate on diverse mathematics benchmarks, including GaoKao2023en (Chinese GaoKao Community, 2024), AIME24 (AIME, 2024), AIME25 (AIME, 2025), AMC (AMC, 2023), MATH-500 (Hendrycks et al., 2021), and OlympiadBench (He et al., 2024). To assess out-of-distribution (OOD) generalization, we evaluate on the scientific reasoning benchmark, GPQA-Diamond (Rein et al., 2024). For all benchmarks, we report pass@1 accuracy via greedy decoding. Vanilla GRPO is trained with our data and hyperparameters, and LUFFY on our data with its original parameters. For Simple-RL and Oat-Zero, we evaluate their publicly available weights.
@@
We train all models for 10 epochs using the verl framework (Sheng et al., 2025), reporting results from the best-performing checkpoint. The maximum response length is 2048 tokens (8192 for the LongCoT model). Consistent with recent studies (Liu et al., 2025b; Yu et al., 2025; Yan et al., 2025), we set the KL divergence penalty to zero to maximize policy exploration. A comprehensive list of hyperparameters is detailed in Appendix C.
@@ 4.2 Main Results
In this section, we present the primary evaluation of Scaf-GRPO, focusing on pass@1 performance across diverse model architectures. The results highlight the method’s performance advantage over Vanilla GRPO and competing prefix-based baselines. Comprehensive supplementary analyses are detailed in Appendix I, covering robustness checks using the avg@16 metric, comparisons against an expanded suite of methods such as DAPO Yu et al. (2025) and DeepScaleR Luo et al. (2025).
@@
To verify that the benefits of Scaf-GRPO are not confined to a single model family, we extend our evaluation to the Llama-3.2-3B-Instruct model (Dubey et al., 2024). As detailed in Table 1, our framework demonstrates strong generalization. While vanilla GRPO provides a significant uplift over the base model, Scaf-GRPO achieves a further relative improvement of 10.3% in average performance. This confirms Scaf-GRPO is a model-agnostic method, capable of enhancing reasoning abilities beyond the Qwen series.
@@ 4.3 Ablation Study
We conduct a series of ablation studies on the Qwen2.5-Math-7B model (see Table 2). Detailed investigations into the guidance exemption period and data filtering strategies are presented in Appendix G.1 and Appendix G.2, respectively.
Necessity and robustness of the guidance exemption period.
To validate the necessity and robustness of the guidance exemption period (Phase 1), we conduct detailed ablation studies in Appendix G.1. Our experiments show that applying scaffolding from the very beginning leads to a 9.2% relative performance drop on Qwen2.5-Math-7B compared to the full framework, confirming that an initial phase of autonomous exploration is crucial to prevent hint dependency. Furthermore, sensitivity analysis reveals that the method is highly stable across exemption durations ranging from 10% to 40%, where the model maintains a high-performance plateau between 49.5% and 50.9%. This validates our selection of 15% as an optimal and robust configuration.
@@
Our methodology is founded on the hypothesis that progressive guidance, from abstract concepts to concrete steps, is superior to simply providing a direct solution. To test this, we evaluate a “Solution-Only” variant that bypasses the hierarchy and immediately provides the most concrete hint, which is $H_{\text{solution}}$. This results in a significant performance degradation of 4.9% compared to the full model. This confirms our hypothesis: compelling the model to first engage with higher-level reasoning fosters more generalizable skills.
@@
Impact of hint quality.
We investigate the correlation between hint quality and student performance using a multi-faceted rubric (accuracy, minimality, clarity, and structural coherence). Employing an LLM-as-a-Judge, we observe that higher-quality hints lead to superior downstream outcomes. Notably, DeepSeek-R1 achieved a higher hint quality score, outperforming Qwen2.5-72B-Instruct. This translated to significant performance gains for the student model (Qwen2.5-Math-7B), yielding a relative accuracy improvement of 4%. We defer the complete evaluation rubric, the specific prompt used for the LLM judge, and detailed experimental results to Appendix H.